### v22

# Simple Lifelong Learning Machines

Abstract
In lifelong learning, data are used to improve performance not only on the present task, but also on past and future (unencountered) tasks. While typical transfer learning algorithms can improve performance on future tasks, their performance on prior tasks degrades upon learning new tasks (called forgetting). Many recent approaches for continual or lifelong learning have attempted to maintain performance on old tasks given new tasks. But striving to avoid forgetting sets the goal unnecessarily low. The goal of lifelong learning should be to use data to improve performance on both future tasks (forward transfer) and past tasks (backward transfer). In this paper, we show that a simple approach—representation ensembling—demonstrates both forward and backward transfer in a variety of simulated and benchmark data scenarios, including tabular, vision (CIFAR-100, 5-dataset, Split Mini-Imagenet, Food1k, and CORe50), and speech (spoken digit), in contrast to various reference algorithms, which typically failed to transfer either forward or backward, or both. Moreover, our proposed approach can flexibly operate with or without a computational budget.
Index Terms:
continual/lifelong learning, forward transfer, backward transfer, replay
I Introduction
Learning is a process by which an intelligent system improves performance on a given task by leveraging data [1]. In classical machine learning, the system is often optimized for a single task [2, 3]. While it is relatively easy to simultaneously optimize for multiple tasks (multi-task learning) [4], it has proven much more difficult to sequentially optimize for multiple tasks [5, 6]. Specifically, classical machine learning systems, and natural extensions thereof, exhibit “catastrophic forgetting” when trained sequentially, meaning their performance on the prior tasks drops precipitously upon training on new tasks [7, 8, 9]. However, learning could be lifelong, with agents continually building on past knowledge and experiences, improving on many tasks given data associated with any task. For example, in humans, learning a second language often improves performance in an individual’s native language [10].
In lifelong learning, where tasks arrive sequentially, the ability to transfer knowledge across tasks is characterized by two complementary objectives: forward transfer and backward transfer. Forward transfer facilitates accelerated learning in new tasks using previous knowledge. In contrast, backward transfer evaluates the impact of new learning on previously encountered tasks. Achieving both positive forward and backward transfer is crucial for an effective lifelong learner. However, as we will demonstrate, many existing lifelong learning algorithms do not enable forward transfer to future tasks, and most do not exhibit positive backward transfer to previously learned tasks.
In this paper, we propose a general and simple approach for lifelong learning which can be used with many existing encoder models. Specifically, we focus our approach on ensembling deep networks (Simple Lifelong Learning Networks, SiLLy-N). Additionally, we demonstrate how the same approach can be generalized for lifelong learning based on ensembling decision forests (Simple Lifelong Learning Forests, SiLLy-F). Table I and Figure 3 show our proposed approaches grow linearly with the task number and can flexibly operate with both growing and constant resources depending on the available computation budget. Moreover, we explore our proposed algorithm as compared to a number of reference algorithms on an extensive suite of numerical experiments that span simulation, vision datasets including CIFAR-100, 5-dataset, Split Mini-Imagenet, and Food1k, as well as the spoken digit dataset. Figure 1 illustrates that our algorithm outperforms all the reference algorithms in terms of forward, backward, and overall transfer on different vision and speech datasets. Ablation studies indicate the degree to which the amount of representation or storage capacity and replaying old task data impact performance of our algorithms. All our code and experiments are open source to facilitate reproducibility.
The subsequent organization of the paper can be summarized as:
  1. 1.
    Section II presents a discussion of relevant algorithms, highlighting their architecture and computational complexity in comparison to SiLLy-N and SiLLy-F.
  2. 2.
    Section III introduces the learning environment and proposes three transfer statistics.
  3. 3.
    Section IV details the design of SiLLy-N and SiLLy-F, incorporating insights from various ensembling approaches.
  4. 4.
    Sections V and VI provide empirical evaluations of SiLLy-N and SiLLy-F on both simulated and benchmark datasets.
  5. 5.
    Section VII concludes with a discussion of strengths, limitations, and directions for future research.
[…]
II Related Works
In this work, we propose a lifelong learning approach based on representation ensembling which can flexibly operate in both resource growing and resource constrained modes. Existing resource constrained algorithms, such as EWC [11], Online EWC [12], EWC++ [13], SI [14], and LwF [15], use regularization techniques to exploit the stability-plasticity trade-off to mitigate forgetting.
On the contrary, the resource growing approaches add resources as they face more tasks. Authors from [16] used a weighted ensemble of learners in a streaming setting with distribution shift. TrAdaBoost [18] boosts ensemble of learners to enable transfer learning. In continual learning scenarios, many algorithms have been built on these ideas by ensembling dependent representations. For example, Learn++ [19] boosts ensembles of weak learners learned over different data sequences in class incremental lifelong learning settings [20]. Model Zoo [21] uses the same boosting approach in task incremental lifelong learning scenarios. Another group of algorithms, ProgNN [22] and DF-CNN [23] learn a new “column” of nodes and edges with each new task, and ensembles the columns for inference (such approaches are commonly called ‘modular’ now). The primary difference between ProgNN and DF-CNN is that ProgNN has forward connections to the current column from all the past columns. This creates the possibility of forward transfer while freezing backward transfer. However, the forward connections in ProgNN render it computationally inefficient for a large number of tasks. DF-CNN gets around this problem by learning a common knowledge base and thereby, creating the possibility of backward transfer.
Recently, many other modular approaches have been proposed in the literature that improve on ProgNN’s capacity growth. These methods consider the capacity for each task being composed of modules that can be shared across tasks and grown as necessary. For example, PackNet [24] starts with a fixed capacity network and trains for additional tasks by freeing up portion of the network capacity using iterative pruning. Veniat et al. [25] trains additional modules with each new task, and the old modules are only used selectively. [26] improved the memory efficiency of the modular methods by adding new modules according to the complexity of the new tasks. Authors in [27] proposed non-parametric factorization of the layer weights that promotes sharing of the weights between tasks. However, all of modular methods described above lack backward transfer because the old modules are not updated with the new tasks. Dynamically Expandable Representation (DER) [28] proposed an improvement over the modular approaches where the model capacity is dynamically expanded and the model is fine-tuned by replaying a portion of the old task data along with the new task data. This approach achieves backward transfer between tasks as reported by the authors in the experiments. Another modular approach, proposed by Kang et al. [29], uses sub-networks inspired by the lottery ticket hypothesis [30] and dynamically expands the network capacity with the number of tasks.
Another strategy for building lifelong learning machines is to use total or partial replay [31, 32, 33, 34, 35]. Replay approaches keep the old data and replay them when faced with new tasks to mitigate catastrophic forgetting. Many recent approaches combine replay with other transfer learning strategies like knowledge distillation [37] to improve storage efficiency of replay data. Recently, hypernetwork based old network weight generation accompanied by synthetic replay [38] demonstrated potential of memory-efficient operation in a learning environment with a huge number of tasks. Another replay algorithm, MER [39], aligns the gradient update between task datasets while replaying old task data. However, as we illustrated in the main text, previously proposed replay algorithms do not demonstrate positive backward transfer in our experiments, though they often do not forget as much as other approaches.
Our approach builds directly on previously proposed modular and replay approaches with one key distinction: in our approach, representations are learned independently. Independent representations also have computational advantages, as doing so merely requires quasilinear time and space, and can be learned in parallel.
III Mathematical Framework
III-A The lifelong learning objective
In a lifelong learning setting, we consider a sequence of tasks $\mathcal{T}=\{1,2,\dots,T\}$, where the tasks are known during training and testing. Each task $t\in\mathcal{T}$ shares the same input space, $\mathcal{X}\subset\mathbb{R}^{D}$, with class labels $\mathcal{Y}=\{1,\dots,K_{t}\}$. The tasks arrive sequentially, and within each task $t$, the dataset $\mathbf{S}^{t}=\{(X_{i},Y_{i})\}_{i=1}^{n_{t}}$ is sampled $iid$ from a fixed distribution $\mathcal{D}_{t}$, with $\sum_{t=1}^{T}n_{t}=n$.
Given access to all observed datasets $\bigcup_{t=1}^{T}\mathbf{S}^{t}$, the goal is to find a learner $f$ that minimizes the overall generalization error across all tasks:
$\begin{array}[]{ll}\textnormal{minimize}&\sum_{t=1}^{T}\mathcal{E}_{f}^{t}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})\\ \textnormal{subject to }&f\in\mathcal{F}\end{array}\enspace,$ (1)
where $\mathcal{E}_{f}^{t}$ is the generalization error or expected risk on task $t$ and the learner will have access to a total of $T$ datasets after $T$ tasks, $\bigcup_{t=1}^{T}\mathbf{S}^{t}$  [^1] More generally, we may have $J$ datasets, where $J\neq T$ and each dataset may be associated with the target distributions of multiple tasks. For simplicity, we do not consider such scenarios further at this time.. Risk is defined as the expected task-specific loss $\ell_{t}:\mathcal{Y}\times\mathcal{Y}\rightarrow[0,\infty)$ in task $t$ (see [40] for a detailed formulation on different out-of-distribution learning scenarios).
III-B Lifelong learning evaluation criteria
Others have previously introduced criteria to evaluate transfer, including forward and backward transfer [41, 42, 43, 25]. Pearl [44] introduced the transfer benefit ratio, which builds directly off relative efficiency from classical statistics [45]. We define three notions of transfer building on relative efficiency.
Definition 1 (Transfer).
Overall transfer of algorithm $f$ for a given Task $t$ is:
$\mathsf{Transfer}^{t}(f):=\log\frac{\mathcal{E}^{t}_{f}(\mathbf{S}^{t})}{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})}.$ (2)
In words, Equation 2 quantifies the extent that a learner $f$, is able to improve the performance on task $t$, when using the data on all other tasks, $(1,\dots,T)$. It does so by computing the ratio of generalization error on task $t$ while only training on data from task $t$, relative to the generalization error on task $t$ when training on the data from all tasks.
Forward transfer quantifies how much performance a learner transfers forward to future tasks, given prior tasks.
Definition 2 (Forward Transfer).
The forward transfer of $f$ for task $t$ is :
$\mathsf{Forward~Transfer}^{t}(f):=\log\frac{\mathcal{E}^{t}_{f}(\mathbf{S}^{t})}{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{t}\mathbf{S}^{t^{\prime}})}.$ (3)
Forward transfer is identical to transfer as defined in Equation 2, except that it only considers the relation between training on data from task $t$ as compared to training on the prior tasks, $(1,\dots,t)$, excluding future tasks, $(t+1,\dots,T)$.
Backward transfer quantifies how much a learner transfers backward to previously observed tasks, in light of new tasks.
Definition 3 (Backward Transfer).
The backward transfer of $f$ for Task $t$ is:
$\mathsf{Backward~Transfer}^{t}(f):=\log\frac{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{t}\mathbf{S}^{t^{\prime}})}{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})}.$ (4)
Backward transfer is also like Equation 2, except it compares the performance when training on all prior tasks, $(1,\dots,t)$, with training on all tasks, $(1,\dots,T)$.
Note that $\mathsf{Transfer}$ can be decomposed into $\mathsf{Forward~Transfer}$ and $\mathsf{Backward~Transfer}$ (see Appendix -A for details):
$\mathsf{Transfer}^{t}(f)=\mathsf{Forward~Transfer}^{t}(f)+\mathsf{Backward~Transfer}^{t}(f).$ (5)
The above equation underscores the sequential progression of tasks in lifelong learning, indicating that the total knowledge a learner acquires is derived from both previously encountered tasks (forward transfer) and those yet to be encountered (backward transfer). While forward transfer is relatively straightforward to achieve [46], backward transfer presents a significant challenge due to the issue of catastrophic forgetting. In particular, a learner that demonstrates strong forward transfer ($\mathsf{Forward~Transfer}^{t}(f)>0$) but experiences negative backward transfer ($\mathsf{Backward~Transfer}^{t}(f)<0$) from all future tasks will eventually lose the learned knowledge, resulting in $\mathsf{Transfer}^{t}(f)\leq 0$ for the task.
Another paper [25], concomitantly introduced transfer and forgetting (backward transfer). Their statistics are the same as ours, except they do not use a $\log$. We opted for a $\log$ to address numerical stability issues in comparing small numbers. Because log is a monotonic function, the order of ranking algorithms is preserved (Appendix Figure 2 shows a version of Figure 1, but using Veniat’s statistics, which is nearly visually identical). By virtue of introducing $\mathsf{Forward~Transfer}$ here, we can identify the inherent trade-off between forward and backward transfer, for a fixed amount of total transfer. Apart from the above statistics, we also report accuracy per task.
Definition 4 (Accuracy).
The accuracy of algorithm $f$ on task $t$ after observing total $T$ datasets is:
$\mathsf{Accuracy}^{t}(f):=1-{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})}.$ (6)
III-C Computational Taxonomy of Lifelong Learners
TABLE I: A computational taxonomy of lifelong learners. We show soft-O notation ($\tilde{\mathcal{O}}(\cdot,\cdot)$) as a function of total training samples, $n=\sum_{t}^{T}n_{t}$, where $n_{t}$ is the number of training samples for the $t^{th}$ task and total task, $T$, as well as the common setting where $n$ is proportional to $T$. Parametric, semi-parametric, and non-parametric algorithms have parameters that remain fixed, grow slowly, and scale proportionally to $n$, respectively.
Parametric Capacity Space Time Examples
[…]
parametric 111 111 111 111 111 SiLLy-N-M, SiLLy-F-M
parametric 111 111 111 $n$ $n$ O-EWC [12], SI [14], LwF [15]
parametric 111 $T$ $n$ $nT$ $n^{2}$ EWC [11]
parametric 111 $n$ $n$ $nT$ $n^{2}$ Total Replay
semi-parametric $T$ $T^{2}$ $n^{2}$ $nT$ $n^{2}$ ProgNN[22]
semi-parametric $T$ $T$ $n$ $n$ $n$ DF-CNN [23]
semi-parametric $T$ $T+n$ $n$ $n$ $n$ SiLLy-N, Model Zoo[21], DER [28], LMC[26]
non-parametric $n$ $n$ $n$ $n$ $n$ SiLLy-F, IBP-WF [27]
In the past 30 years, a number of lifelong learning algorithms have attempted to overcome catastrophic forgetting [47, 48]. These algorithms can be broadly classified into three categories: parametric, semi-parametric, and non-parametric approaches, based on their representational capacity and how it scales with the number of tasks and data samples. Table I shows a computational taxonomy of lifelong learners as a function of the sample size $n$ and the total task $T$, as well as the common scenario where the sample size is fixed per task and therefore proportional to the number of tasks, $n\propto T$. For time complexity, we only consider the training time complexity. The space complexity of the learner refers to the amount of memory space needed to store the learner [49]. We also study the representation capacity of these algorithms. Capacity is defined as the size of the subset of hypotheses that is achievable by the learning algorithm [50]. For a detailed discussion and assumptions behind the complexity analysis, see Appendix A-C.
IV Representation ensembling algorithms
Figure 2: Schemas of composable hypotheses. A. Single task learner. B. Ensembling decisions (as output by the channels) is a well-established practice, including random forests and gradient boosted trees. C. Learning a joint representation or D. learning future representations depending on the past encoders was previously used in lifelong learning scenarios, but encoders were not trained independently as in E. Note that the new encoders in E interact with the previous encoders through the channel layer (indicated by red arrows), thereby, enabling backward transfer. Again the old encoders interact with the future encoders (indicated by black arrows), thereby, enabling forward transfer.
Shannon proposed that a learned hypothesis can be decomposed into three components: an encoder, a channel, and a decoder [51, 52]: $h(\cdot)=w\circ v\circ u(\cdot)$. Figure 2 shows these three components as the building blocks of different learning schemas. The encoder, $u:\mathcal{X}\mapsto\tilde{\mathcal{X}}$, maps an $\mathcal{X}$-valued input into an internal representation space $\tilde{\mathcal{X}}$ [53, 54]. The channel $v:\tilde{\mathcal{X}}\mapsto\Delta_{\mathcal{Y}}$ maps the transformed data into a posterior distribution (or, more generally, a score). Finally, a decoder $w:\Delta_{\mathcal{Y}}\mapsto\mathcal{Y}$, produces a predicted label.
A canonical example of a single learner depicted in Figure 2A is a decision tree. Importantly, one can subsample the training data to learn different components of the tree [55, 56, 57]. For example, one can use a portion of data to learn the tree structure (which is the encoder). Then, by pushing the remaining data (sometimes called the ‘out-of-bag’ data) through the tree, one can learn posteriors in each leaf node (which are the channel). The channel thus gives scores for each data point denoting the probability of that data point belonging to a specific class. Using separate sets of data to learn the encoder and the channel results in less bias in the estimated posterior in the channels as in ‘honest trees’ [55, 56, 57]. Finally, the decoder provides the predicted class label using $\operatorname*{arg\,max}$ over the posteriors from the channel.
One can generalize the above decomposition by allowing for multiple encoders, as shown in Figure 2B. Given $B$ different encoders, one can attach a single channel to each encoder, yielding $B$ different channels. Doing so requires generalizing the definition of a decoder so that it would operate on multiple channels. Such a decoder ensembles the decisions, because here each channel provides the final output based on the encoder. This is the learning paradigm behind bagging [58] and boosting [59]; indeed, decision forests are a canonical example of a decision function operating on an ensemble of $B$ outputs [60].
Although the task specific structure in Figure 2B can provide useful decision on the corresponding task, they cannot, in general, provide meaningful decisions on other tasks, because those tasks might have completely different class labels. Therefore, in the multi-head structure (Figure 2C) a single encoder is used to learn a joint representation from all the tasks, and a separate channel is learned for each task to get the score or class conditional posteriors for each task, which is followed by each task specific decider [11, 12, 14].
Modular approaches, such as ProgNN and LMC (Figure 2D), have both multiple encoders and decoders. Connections from past to future encoders enables forward transfer. However, they freeze backward transfer.
Our approach also uses multiple encoders and decoders (Figure 2E). Unlike modular approaches, we allow interaction among encoders through the channels, including both forward and backward interactions. The result is that the channels ensemble representations (learned by the encoders), rather than decisions (learned by the channels as in Figure 2 B). In our algorithms, we push all the data through each encoder, and each channel learns and ensembles across all encoders. When each encoder has learned complementary representations, the channels can leverage that information to improve over single task performance. This approach has applications in few-shot and multiple task scenarios, as well as lifelong learning.
IV-A Our representation ensembling algorithms
Figure 3: Space and time complexity as a function on number of tasks in CIFAR 10X10. Section VI describes the experimental setup and architecture of the convolutional network used. Top left: Model size, SiLLy can flexibly operate between resource constrained and growing modes. Top right: Memory consumed by SiLLy-N is dominated by the encoder size. Bottom left: SiLLy takes less training time compared to other baselines. Bottom right: Inference time taken by different lifelong learners for 100010001000 samples. Bottom row has log y-axis.
Figure 2E shows a general structure of our algorithm. As data from a new task arrives, the algorithm first builds a new encoder. Then, it builds the channel for this new task by pushing the new task data through all existing encoders. Thus the channel integrates information across all existing encoders using the new task data, thereby enabling forward transfer. At the same time, if it stores old task data (or can generate such data), it can push that data through the new encoders to update the channels from the old tasks, thereby enabling backward transfer. In either case, new test data are passed through all existing encoders and corresponding channels to make a prediction (see appendix for detailed description of this approach).
IV-A1 Simple Lifelong Learning Networks in resource growing mode
A Simple Lifelong Learning Network (SiLLy-N) ensembles deep networks. For each task, the encoder $u_{t}$ in SiLLy-N is the “backbone” of a deep network (DN). Thus, each $u_{t}$ maps an element of $\mathcal{X}$ to an element of $\mathbb{R}^{d}$, where $d$ is the number of neurons in the penultimate layer of DN. The channels are learned by averaging the outputs from decision forests [61, 60] trained on the $d$ dimensional representations of $\mathcal{X}$. See Appendix Figure 3 where we conduct experiments on a CIFAR 10X10 (described later in Section VI-B1) while varying the number of decision trees per channel. Appendix Figure 3 shows that decision forest-based channels are robust to hyperparameter perturbation. Other algorithms could also be used to learn the channels, though we do not pursue them here. The decoder $w_{t}$ outputs the $\operatorname*{arg\,max}$ to produce a single prediction.
IV-A2 Simple Lifelong Learning Networks in resource constrained mode
The above resource growing approach is ideal when the upcoming tasks become more and more complex and there is no constraint imposed by the computation and storage budget available. However, real-world scenarios often impose computational constraints. In the constant resource mode, we stop building new encoders after we have reached the computation and the storage budget imposed by the user. As new tasks arrive, we only learn new channels associated with new tasks using the old encoders. Note that this approach completely excludes the need to save old task data after we have reached the budget.
Hereafter, we will use suffix ‘-M’ after the algorithm name whenever we use resource constrained operation of SiLLy-N. Here $M$ is the total number of encoder allowed by the budget.
IV-A3 Additional realization of our approach using random forest as encoder
Simple Lifelong Learning Forest (SiLLy-F) ensembles decision trees or forests. For each task, the encoder $u_{t}$ of SiLLy-F is the representation learned by a decision forest. The channel then learns the class-conditional posteriors by populating the forest leaves with out-of-task samples, as in “honest trees” [55, 56, 57]. Each channel outputs the posteriors averaged across the collection of forests learned over different tasks. The decoder $w_{t}$ outputs the argmax to produce a single prediction.
Note that the amount of additional representation capacity added per task by SiLLy-F is a function of the amount and complexity of the data for a new task. Contrast this with SiLLy-N  and other deep net based modular or representation ensembling approaches, which a priori choose how much additional representation to add, prior to seeing all the new task data. So, SiLLy-F has capacity, space complexity, and time complexity scale with the complexity and sample size of each task. In contrast, ProgNN, SiLLy-N (and others like it) have a fixed capacity for each task, even if the tasks have very different sample sizes and complexities.
Figure 3 top left shows the model size for our proposed approach grows linearly with the number of tasks. Moreover, the memory consumed by the new channels is negligible compared to the memory required to store the encoders (Figure 3 top right). The time required for inference on 100010001000 testing points (Figure 3 bottom right) is an order of magnitude lower compared to the time required to train a new encoder with 500500500 samples(Figure 3 bottom left).
V Simulation Data Study
Figure 4: SiLLy-N demonstrates forward and backward transfer. (A) 750 samples from: (Ai) Gaussian XOR, (Aii) XNOR, which has the same optimal discriminant boundary as XOR, and (Aiii) R-XOR, which has a discriminant boundary that is uninformative, and therefore adversarial, to XOR. (Bi) Generalization error for XOR, and (Bii) XNOR. SiLLy-N outperforms DN on XOR when XNOR data is available, and on XNOR when XOR data are available. (Biii) Forward and backward transfer of SiLLy-N are positive for all sample sizes. (Ci) In an adversarial task setting, SiLLy-N gracefully forgets XOR, whereas DN  catastrophically forget and interfere. (Cii) $\mathsf{Backward~Transfer}$ is maximum positive with respect to XOR when the optimal decision boundary of $\theta$-XOR is similar to that of XOR (e.g. angles far from $45^{\circ}$), and negative otherwise. The dashed line shows the regression line fitted through the original points. (Ciii) $\mathsf{Backward~Transfer}$ is a nonlinear function of the source training sample size (XOR sample size is fixed at 500500500).
V-A Forward and backward transfer in a simple environment
Consider a very simple two-task environment: Gaussian XOR and Gaussian Exclusive NOR (XNOR) (Figure 4A, see Appendix A-D for details). The two tasks share the exact same discriminant boundaries: the coordinate axes. Thus, transferring from one task to the other merely requires learning a bit flip of the class labels. We sample a total 750 samples from XOR, followed by another 750 samples from XNOR.
SiLLy-N and deep network (DN) achieve the same generalization error on XOR when training with XOR data (Figure 4Bi). But because DN does not account for a change in task, when XNOR data appear, DN performance on XOR deteriorates (it catastrophically forgets). In contrast, SiLLy-N continues to improve on XOR given XNOR data, demonstrating backward transfer. Now consider the generalization error on XNOR (Figure 4Bii). Both SiLLy-N and DN are at chance levels for XNOR when only XOR data are available. When XNOR data are available, DN must unlearn everything it learned from the XOR data, and thus its performance on XNOR starts out nearly maximally inaccurate, and quickly improves. On the other hand, because SiLLy-N can leverage the encoder learned using the XOR data, upon getting any XNOR data, it immediately performs quite well, and then continues to improve with further XNOR data, demonstrating forward transfer (Figure 4Biii). SiLLy-N demonstrates positive forward and backward transfer for all sample sizes, whereas DN fails to demonstrate neither forward nor backward transfer, and eventually catastrophically forgets the previous tasks.
V-B Forward and backward transfer for adversarial tasks
In the context of lifelong learning, we informally define a task $t$ to be adversarial with respect to task $t^{\prime}$ if the true joint distribution of task $t$, without any domain adaptation, impedes performance on task $t^{\prime}$. In other words, training data from task $t$ can only add noise, rather than signal, for task $t^{\prime}$. An adversarial task for Gaussian XOR is Gaussian XOR rotated by $45^{\circ}$ (R-XOR) (Figure 4Aiii). Training on R-XOR therefore impedes the performance of SiLLy-N  on XOR, and thus backward transfer becomes negative, demonstrating graceful forgetting [62] (Figure 4Ci).
To further investigate this relationship, we design a suite of R-XOR examples, generalizing R-XOR from only $45^{\circ}$ to any rotation angle between $0^{\circ}$ and $90^{\circ}$, sampling 100 points from XOR, and another 100 from each R-XOR (Figure 4Cii). Note that we could not run the experiment for a lot of Monte Carlo repetition to have a smooth curve and hence we have shown a regressed curve fitted to the low repetition noisy curve. As the angle increases from $0^{\circ}$ to $45^{\circ}$, $\mathsf{Backward~Transfer}$ gradually decreases for SiLLy-N. The $45^{\circ}$-XOR is the maximally adversarial R-XOR. Thus, as the angle further increases, $\mathsf{Backward~Transfer}$ increases back up to 0.18$\approx$ 0.18≈ 0.18 at $90^{\circ}$, which has an identical discriminant boundary to XOR. Moreover, when $\theta$ is fixed at $25^{\circ}$, $\mathsf{Backward~Transfer}$ increases at different rates for different sample sizes of the source task (Figure 4Ciii).
Together, these experiments indicate that the amount of transfer can be a complicated function of (i) the difficulty of learning good representations for each task, (ii) the relationship between the two tasks, and (iii) the sample size of each.
VI Benchmark data Study
For benchmark data, we build SiLLy-N encoders using the network architecture described in [31]. We use the same network architecture for all benchmarking models. For the following experiments, we consider two modalities of real data: vision and language.
VI-A Reference algorithms
We compared our approaches to 161616 reference lifelong learning methods. Among them five are resource growing as well as modular approach: ProgNN [22], DF-CNN [23], LMC [26], Model Zoo[21], CoSCL [63]. Note that ”Model Zoo” was published after our work was archived on arXiv, and the authors have built on our work (personal communications). Other reference algorithms are resource constrained: Elastic Weight Consolidation (EWC) [11], Online-EWC (O-EWC) [12], Synaptic Intelligence (SI) [14], Learning without Forgetting (LwF) [15], RanDumb [64] and “None”.
We also compare two variants of exact replay (Total Replay and Partial Replay) using the code provided in [31]. Both Total and Partial Replay store all the data they have ever seen, but Total Replay replays all of it upon acquiring a new task, whereas Partial Replay replays $N$ samples, randomly sampled from the entire corpus, whenever we acquire a new task with $N$ samples. The above two replay approaches can be considered as two variants of GDumb [65]. Additionally, we have compared our approach with more constrained ways of replaying old task data, including Averaged Gradient Episodic Memory (A-GEM) [66], Experience Replay (ER) [34] and Task-based Accumulated Gradients (TAG) [67].
For the baseline “None”, the network was incrementally trained on all tasks in the standard way while always only using the data from the current task. The implementations for all of the algorithms are adapted from open source codes [23, 68]; for implementation details, see Appendix A-B.
Figure 5: Ablation experiments on SiLLy-N using CIFAR 10X10. Top row: SiLLy-N uniquely shows both positive forward and backward transfer while operating with the same number of parameters as other baseline constant parameter approaches. Bottom row: Replaying old task data impacts backward transfer while keeping forward transfer unchanged for SiLLy-N.
Figure 6: Pretrained encoders on CIFAR 10X10. Using pretrained encoders results in better forward transfer and accuracy for SiLLy-N.
VI-B Core benchmarks
VI-B1 CIFAR 10X10
The CIFAR 100 challenge [69], consists of 50,000 training and 10,000 test samples, each a 32x32 RGB image of a common object, from one of 100 possible classes, such as apples and bicycles. CIFAR 10x10 divides these data into 10 tasks, each with 10 classes [23] (see Appendix A-E for details).
SiLLy-N and Model Zoo demonstrate positive forward and backward transfer for every task in CIFAR 10x10, in contrast, other algorithms do not exhibit any positive backward transfer (Figure 1 first column). Moreover, they retained their accuracy while improving transfer (Figure 1 first column, first and fourth rows). ProgNN had a similar degree of forward transfer, but zero backward transfer, and requires quadratic space and time in sample size, unlike SiLLy-N which requires quasilinear space and time.
VI-B2 Spoken Digit
In this experiment, we used the Spoken Digit dataset [70]. As shown in Figure 1 second column, SiLLy-N shows positive backward and forward transfer between the spoken digit tasks, in contrast to other methods, some of which show only forward transfer, others show only backward transfer, with none showing both and some showing neither. See Appendix A-E for details of the experiment.
VI-B3 FOOD1k 50X20 Dataset
In this experiment, we use Food1k which is a large scale vision dataset consisting of 100010001000 food categories from Food2k [71]. FOOD1k 50X20 splits these data into 505050 tasks with 202020 classes each. For each class, we randomly sampled 606060 samples per class for training the models and used rest of the data for testing purpose. Because on the CIFAR experiments Model Zoo performs the best among the reference resource growing models, and LwF is the best performing resource constrained algorithm, we only use them as the reference models for the large scale experiment to avoid heavy computational cost. As shown in Figure 1 third column, SiLLy-N performs the best among all the algorithms on this large dataset.
See Appendix A-E for experiments with datasets having more samples per task. In lifelong learning, we are often primarily concerned with situations in which we have a small number of samples per task. If we have enough samples per task, the learning agent does not need to transfer knowledge from other tasks. However, below we also experiment with non-trivial lifelong learning setting where sample per task is high.
These experiments indicate that SiLLy-N has positive transfer in learning environments with various classes and sample sizes.
VI-C Ablation experiments
Our proposed algorithms can improve performance on all the tasks (past and future) by both growing additional resources and replaying data from the past tasks. Below we do two ablation experiments using CIFAR 10X10 to measure the relative contribution of resource growth and replay to the performance of our proposed algorithms.
Constrained resource experiment
In this experiment, we ablate the capability of SiLLy-N to grow additional resources after learning 444 encoders. We also reduce the number of channels and nodes at each encoder layer by four times to keep the total number of parameters similar to the other constant-resource-algorithms. As shown in the top row of Figure 5, SiLLy-N-4 still shows positive forward and backward transfer with constant resources. However, the accuracy for SiLLy-N-4 gets reduced compared to that of resource growing SiLLy-N in Figure 1. Note that all the baseline algorithms have negative backward transfer. This experiment indicates that constant resource mode operation for SiLLy-N may be advantageous when we have a lot of tasks to learn and have a decent amount of storage budget available. We will elaborate the above point later with a large scale dataset (food1k).
Controlled replay experiment
[…]
These experiments indicate that (i) constraining the resource growth results in lower accuracy while still achieves positive forward and backward transfer for the algorithm, (ii) lowering the amount of replay lowers backward transfer without any effect on forward transfer.
Experiment using pretrained encoders
We explore the effect of using pretrained encoders on the performance of SiLLy-N. For this experiment only, we use ResNet 50 and vision transformer ViT_B16 (provided in keras-vit package) [72]. We freeze all the layers excluding the final two linear layers during training. Pretraining the encoders results in better accuracy and forward transfer, but less backward transfer for SiLLy-N (Figure 6). Vision transformers achieved lower accuracy, as expected with the small samples sizes used here [73]. Pretrained network requires less training epochs for each task (with early stopping). In this experiment, one forward pass through ViT takes 5$\sim$ 5∼ 5 seconds and training a ViT encoder (202020 epochs) takes 4$\sim$ 4∼ 4 minutes using an Apple M1 Max chip and 646464 GB of RAM.
VI-D Adversarial analysis
Figure 7: Adversarial CIFAR 10x10 experiments. Left: Permuting the task sequenc does not affect $\mathsf{Transfer}$ for SiLLy-N. The mean position, indicated by the black dash, and the spread of the dots on both sides of the mean remain nearly the same. Middle: Shuffling class labels within tasks two through nine with 500 samples each demonstrates SiLLy-N can still achieve positive backward transfer, and that the other algorithms still fail to transfer. Right: SiLLy-N is nearly invariant to rotations, whereas other approaches are more sensitive to rotation.
Consider the same CIFAR 10x10 experiment setup in Section VI-B1. In the following experiments, we modify the above setup in different adversarial settings.
Shuffled task order experiment
We repeat the same experiment mentioned above 555 times more by permuting the task order. Figure 7 left column shows the mean as well as the spread of $\mathsf{Transfer}$ over all tasks remain the same for different shuffled task orders. Note that the encoders and thus all the channels remain the same regardless of the above task order permutation after all the tasks have been introduced, hence the distribution of $\mathsf{Transfer}$ for all the tasks after all the tasks have been introduced remain similar.
Label shuffle experiment
In this experiment, for Task 2 through 10, randomly permute the class labels within each task, rendering each of those tasks adversarial with respect to the first task (because the labels are uninformative). Figure 7 middle column indicates that SiLLy-N show positive backward transfer even with such label shuffling (the other algorithms, except Model Zoo, did not demonstrate positive backward transfer).
Rotation experiment
Consider a Rotated CIFAR experiment, which uses only data from the first task, divided into two subsets of equal size (making two tasks), where the second subset is rotated by different amounts (Figure 7 right column). Backward transfer of SiLLy-N is nearly invariant to rotation angle, whereas the other approaches are much more sensitive to the rotation angle. Note that the zero rotation angle corresponds to the two tasks having identical distributions. The fact that other algorithms fail to transfer even in this setting suggests that they may never be able to achieve a positive backward transfer. See Appendix A-E for an additional experiment using CIFAR 10X10.
These adversarial experiments indicate that SiLLy-N is robust to adversarial perturbations in the source tasks, while most of the other algorithms are not.
VI-E Constant Resource Mode Operation
Figure 8: SiLLy-N in constant resource mode operation. Improvement in Transfer between tasks becomes negligible at nearly 303030 encoders. The user can choose to operate in lower budget with less transfer.
VI-E1 FOOD1k 50X20
The binary distinction we made above, algorithms either build resources or reallocate them, is a false dichotomy, and biologically unnatural. In biological learning, systems develop from building to fixed resources, as they grow from juveniles to adults. To explore this continuum of amount of resources to grow, we experiment on FOOD1k 50X20 dataset using the constant resource mode operation of SiLLy-N as described in Section IV. We evaluate the performance of SiLLy-N for different number of encoder budget. Performance of SiLLy-N saturates after 303030 encoders, though with only 555 encoders, still demonstrates forward and backward transfer (Figure 8).
VI-E2 CORe50 110X5
[…]
VI-F SiLLy-F on tabular data
Figure 9: Our proposed approach can be used with random forest as encoders on tabular data (SiLLy-F). SiLLy-F shows more positive forward and backward transfer while operating with less parameters compared to other baseline approaches on tabular data.
In this experiment, we experiment with SiLLy-F, an additional realization of our approach using random forests as encoders (described in Section IV). We flatten the CIFAR 10X10 data and use them as tabular data. We train two other best performing baseline algorithms, SiLLy-N and Model Zoo and use three fully connected hidden layers, each having 200020002000 nodes, as encoders. As shown in Figure 9, SiLLy-F performs the best among all the approaches. This experiment shows our approach can be used as a general structure to do lifelong learning using other machine learning models as encoder.
VII Discussion
We introduced representation ensembling as a simple approach for lifelong learning. Two specific algorithms, SiLLy-N and SiLLy-F, achieve both forward and backward transfer, by leveraging resources learned for other tasks without undue computational burdens. Our work is well suited for federated learning scenarios, where each data center independently trains a model on its private data and shares only the encoders with other centers [76]. In this paper, we have mainly focused on task-aware setting, because it is simpler. Future work will extend our approach to more challenging task-unaware settings [75]. Our code, including the code for reproducing the experiments in this manuscript, is available from http://proglearn.neurodata.io/.
Acknowledgments
The authors thank the support of the NSF-Simons Research Collaborations on the Mathematical and Scientific Foundations of Deep Learning (NSF grant 2031985). We also thank Raman Arora, Dinesh Jayaraman, Rene Vidal, Jeremias Sulam, Guillermo Sapiro, Michael Powell, and Weiwei Yang for helpful discussions. This work is graciously supported by the Defense Advanced Research Projects Agency (DARPA) Lifelong Learning Machines program through contracts FA8650-18-2-7834 and HR0011-18-2-0025. Research was partially supported by funding from Microsoft Research and the Kavli Neuroscience Discovery Institute.
Co-first Author Contributions
J.D.: Conceptualization, Methodology, Software, Data Curation, Formal Analysis, Experimentation, Theoretical proof (see Appendix D), Writing &\&& Editing; J.T.V.: Conceptualization, Methodology, Supervision, Writing—Review &\&& Editing, and Funding Acquisition.
Appendix A Evaluation Criteria
A-A Representation Ensembling Algorithms
A-A1 Model Architecture
In this paper, we proposed two representation ensembling algorithms, Simple Lifelong Learning Networks (SiLLy-N) and Simple Lifelong Learning Forests (SiLLy-F). The two algorithms differ in their details of how to update encoders and channels, but abstracting a level up they are both special cases of the same procedure. Let SiLLy-X refer to any possible representation algorithm. Algorithms  123, and  4 provide pseudocode for adding encoders, updating channels, and doing inference for any SiLLy-X algorithm.
A-A2 Data Preparation
Whenever the learner gets access to a new task data, we use Algorithm 1 to train a new encoder for the corresponding task. We split the data into two portions — in-task set and held out or out-of-bag set.
A-A3 Training Procedures
In-task set is used to learn the encoder and the indices of the out-of-bag (OOB) data which is returned by Algorithm 1 to be used by Algorithm 2 to learn the channel for the corresponding task. Note that we push the OOB data through the in-task encoder and the whole dataset through the cross-task encoders to update the channel, i.e, learn the posteriors according to the new encoder (see Algorithm 3). Finally, Algorithm 4 does inference on a new sample. Given the task identity, we use the corresponding channel to get the average estimated posterior and predict the class label as the $\operatorname*{arg\,max}$ of the estimated posteriors.
Algorithm 1 Add a new SiLLy-X encoder for a task. OOB = out-of-bag.
1:
2:(1) $t$triangleright current task number
3:(2) $\mathcal{D}_{n}^{t}=(\mathbf{x}^{t},\mathbf{y}^{t})\in\mathbb{R}^{n\times p}\times\{1,\ldots,K\}^{n}$triangleright training data for task $t$
4:
5:(1) $u_{t}$triangleright an encoder trained on task $t$
6:(2) $\mathcal{I}_{OOB}^{t}$triangleright a set of the indices of OOB data
[…]
9:  return $u_{t},\mathcal{I}_{OOB}^{t}$
10:end function
Algorithm 2 Add a new SiLLy-X channel for the current task.
1:
2:(1) $t$triangleright current task number
3:(2) $\mathcal{U}=\{u_{t^{\prime}}\}_{t^{\prime}=1}^{t}$triangleright the set of encoders
4:(3) $\mathcal{D}_{n}^{t}=(\mathbf{x}_{t},\mathbf{y}_{t})\in\mathbb{R}^{n\times p}\times\{1,\ldots,K\}^{n}$triangleright training data for task $t$
5:(4) $\mathcal{I}_{OOB}^{t}$triangleright a set of the indices of OOB data for the current task
[…]
9:  for $t^{\prime}=1,\ldots,t-1$ do
[…]
11:  end for
12:  return $v_{t}$
13:end function
Algorithm 3 Update SiLLy-X channel for the previous tasks.
1:
2:(1) $t$triangleright current task number
3:(2) $u_{t}$triangleright encoder for the current task
4:(3) $\mathcal{D}=\{\mathcal{D}^{t^{\prime}}\}_{t^{\prime}=1}^{t-1}$triangleright training data for old tasks
5:(4) $\mathcal{V}=\{\boldsymbol{v}_{t^{\prime}}\}_{t^{\prime}=1}^{t-1}$triangleright set of all previous task voters
[…]
8:  for $t^{\prime}=1,\ldots,t-1$ do
[…]
10:  end for
11:  return $\mathcal{V}$
12:end function
Algorithm 4 Predicting a class label using SynX.
1:
2:(1) $\boldsymbol{x}\in\mathbb{R}^{p}$triangleright test datum
3:(2) $t$triangleright task identity associated with $\boldsymbol{x}$
4:(3) $\mathcal{U}$triangleright set of all $t$ encoders
5:(4) $v_{t}$triangleright channel for task $t$
[…]
8:  for $t^{\prime}=1,\ldots,t$ do $\triangleright$ get the output $\tilde{\boldsymbol{x}}=\{\tilde{\boldsymbol{x}}_{t^{\prime}}\}_{t^{\prime}=1}^{t}$ from all the encoders
9:   $\tilde{\boldsymbol{x}}_{t^{\prime}}\leftarrow u_{t^{\prime}}.encode(\boldsymbol{x})$
10:  end for
11:  $\hat{\mathbf{p}}\leftarrow v_{t}.predict\_proba(\tilde{\boldsymbol{x}})$triangleright $\hat{\mathbf{p}}$ is a $K_{t}$-dimensional posterior vector
12:  $\hat{y}=\operatorname*{arg\,max}(\hat{\mathbf{p}})$triangleright find the index of the elements in the vector $\hat{\mathbf{p}}$ with maximum value
13:  return $\hat{y}$
14:end function
Figure 1: Error bars (interquartile range, i.e., IQR) for each dot in Figure 1 on different vision and speech benchmark datasets. Error bars for different performance statistics is negligible in comparison with the median performance shown in the main text Figure 1.
Figure 2: Performance summary on vision and audition benchmark datasets using Veniat’s [25]’s statistics. See Figure 1 for caption details. Note that the results here look nearly identical other than the y-axis labels.
Figure 3: Performance of SiLLy-N on CIFAR 10X10 remains nearly unchanged for different number of decision trees per channel.
TABLE I: Hyperparameters for SiLLy-N in CIFAR 10X10, Five Datasets, Split Mini-Imagenet, FOOD1k experiments. Note that we use the same hyperparameters for all the experiments.
Hyperparameters Value
optimizer Adam
learning rate $3\times 10^{-4}$
max_samples (OOB split) 0.670.670.67
n_estimators (decision forest channel) 202020
TABLE II: Hyperparameters for SiLLy-F in tabular CIFAR 10X10 experiments.
Hyperparameters Value
n_estimators 101010
max_depth 303030
max_samples (OOB split) 0.670.670.67
min_samples_leaf 111
A-B Reference Algorithm Implementation Details
The same network architecture was used for all baseline deep learning methods. Following the work in  [31], the ‘base network architecture’ consisted of five convolutional layers followed by two-fully connected layers each containing 2000 nodes with ReLU non-linearities and a softmax output layer. The convolutional layers had 161616, 323232, 646464, 128128128 and 254254254 channels, they used batch-norm and a ReLU non-linearity, they had a 3x3 kernel, a padding of 1 and a stride of 2 (except the first layer, which had a stride of 1). This architecture was used with a multi-headed output layer (i.e., a different output layer for each task) for all algorithms using a fixed-size network. For ProgNN and DF-CNN the same architecture was used for each column introduced for each new task, and in our SiLLy-N this architecture was used for the transformers $u_{t}$ (see above). In these implementations, ProgNN and DF-CNN have the same architecture for each column introduced for each task. Among the reference algorithms, EWC, O-EWC, LwF, SI, Total Replay and Partial Replay results were produced using the repository https://github.com/GMvandeVen/progressive-learning-pytorch. For ProgNN and DF-CNN  we used the code provided in https://github.com/Lifelong-ML/DF-CNN. For all other reference algorithms, we modified the code provided by the authors to match the deep net architecture as mentioned above and used the default hyperparameters provided in the code.
A-C Training Time Complexity Analysis
[…]
  1. 1.
    Each task has the same number of training samples.
  2. 2.
    Capacity grows linearly with the number of trainable parameters in the model.
  3. 3.
    The number of epochs is fixed for each task.
  4. 4.
    For the algorithms with dynamically expanding capacity, we assume the worst case scenario where an equal amount of capacity is added to the hypothesis with an additional task.
Assumption 333 enables us to write time complexity as a function of the sample size. Table I summarizes the capacity, space and time complexity of several reference algorithms, as well as our SiLLy-N and SiLLy-F.
Lifelong learning methods are parametric if they have a representational capacity which is invariant to sample size and task number. Although the space complexity of some of these algorithms grow (because the number of the constraints stored by the algorithms grows, or they continue to store more data), their capacity is fixed. Thus, given a sufficiently large number of tasks with increasing complexity, in general, eventually all parametric methods will catastrophically forget. EWC [11], Online EWC [12], SI [14], and LwF [15] are all examples of parametric lifelong learning algorithms. Our fixed resource algorithms are also parametric.
Lifelong learning methods are semi-parametric if they have a representational capacity which grows slower than sample size. For example, if $T$ is increasing slower than $n$ (e.g., $T\propto\log n$), then algorithms whose capacity is proportional to $T$ are semi-parametric. ProgNN [22] is semi-parametric, nonetheless, its space complexity is $\tilde{\mathcal{O}}(T^{2})$ due to the lateral connections. Moreover, the time complexity for ProgNN also scales quadratically with $n$ when $n\propto T$. Thus, an algorithm that literally stores all the data it has ever seen, and retrains a fixed size network on all those data with the arrival of each new task, would have smaller space complexity and the same time complexity as ProgNN. DF-CNN [23] improves upon ProgNN by introducing a “knowledge base” with lateral connections to each new column, thereby avoiding all pairwise connections. Because these semi-parametric methods have a fixed representational capacity per task, they will either lack the representation capacity to perform well given sufficiently complex tasks, and/or will waste resources for very simple tasks.
Lifelong learning methods are non-parametric if they have a representational capacity which grow in proportion to the number of tasks or data samples. Table I shows the Indian Buffet Process for Weight Factors (IBP-WF) is a notable non-parametric approach alongside SiLLy-F.
Our proposed approaches, SiLLy-N and SiLLy-F, as we will discuss in details in Section IV, eliminate lateral connections between the columns of the network, thus reducing the complexity of the space to $\tilde{\mathcal{O}}(T)$. Moreover, our proposed approaches can adapt flexibly to any of the three categories based on the constraints of the application environment, as illustrated in Table I.
A-C1 Complexity analysis
Consider a lifelong learning environment with $T$ tasks each with $n^{\prime}$ samples, i.e., total training samples, $n=n^{\prime}T$. For all the algorithm with time complexity $\tilde{\mathcal{O}}(n)$, the training time grows linearly with more training samples. We discuss all other algorithms with non-linear time complexity below.
EWC
Consider the time required to train the weights for each task in EWC is $k_{c}n^{\prime}$ and each task adds additional $k_{l}n^{\prime}$ time from the regularization term. Here, $k_{c}$ and $k_{l}$ are both constants. Therefore, time required to learn all the $T$ tasks can be written as:
$k_{c}n^{\prime}+(k_{c}n^{\prime}+k_{l}n^{\prime})+\cdots+(k_{c}n^{\prime}+(T-1)k_{l}n^{\prime})$
$=k_{c}n^{\prime}T+k_{l}n^{\prime}\sum_{t=1}^{T-1}t$
$=k_{c}n^{\prime}T+k_{l}n^{\prime}\frac{T(T-1)}{2}$
$=k_{c}n+0.5k_{l}nT-0.5k_{l}n$
$=\tilde{\mathcal{O}}(nT).$ (11)
Total Replay
Consider the time to train the model on $n^{\prime}$ samples is $k_{c}n^{\prime}$. Therefore, time required to learn all the $T$ tasks can be written as:
$k_{c}n^{\prime}+k_{c}(n^{\prime}+n^{\prime})+\cdots+k_{c}n^{\prime}T$
$=k_{c}n^{\prime}\sum_{t=1}^{T}t$
$=k_{c}n^{\prime}\frac{T(T+1)}{2}$
$=0.5k_{c}nT+0.5k_{c}n$
$=\tilde{\mathcal{O}}(nT)$ (12)
ProgNN
Consider the time required to train each column in ProgNN is $k_{c}n^{\prime}$ and each lateral connection can be learned with time $k_{l}n^{\prime}$. Therefore, time required to learn all the $T$ tasks can be written as:
$k_{c}n^{\prime}+(k_{c}n^{\prime}+k_{l}n^{\prime})+\cdots+(k_{c}n^{\prime}+(T-1)k_{l}n^{\prime})$
$=k_{c}n^{\prime}T+k_{l}n^{\prime}\sum_{t=1}^{T-1}t$
$=k_{c}n^{\prime}T+k_{l}n^{\prime}\frac{T(T-1)}{2}$
$=k_{c}n+0.5k_{l}nT-0.5k_{l}n$
$=\tilde{\mathcal{O}}(nT)$ (13)
A-D Simulation Experiment Details
In each simulation, we constructed an environment with two tasks. For each, we sample 750 times from the first task, followed by 750 times from the second task. These 1,500 samples comprise the training data. We sample another 1,000 hold out samples to evaluate the algorithms. For SiLLy-N, we have used a deep network (DN) architecture with two hidden layers each having 101010 nodes. Similarly, for SiLLy-N experiments we did 100100100 repetitions and reported the results after smoothing it using moving average with a window size of 555.
Gaussian XOR is two class classification problem with equal class priors. Conditioned on being in class 0, a sample is drawn from a mixture of two Gaussians with means ±$\pm\begin{bmatrix}0.5,&0.5\end{bmatrix}^{\mathsf{T}}$, and variances proportional to the identity matrix. Conditioned on being in class 1, a sample is drawn from a mixture of two Gaussians with means ±$\pm\begin{bmatrix}0.5,&-0.5\end{bmatrix}^{\mathsf{T}}$, and variances proportional to the identity matrix. Gaussian XNOR is the same distribution as Gaussian XOR with the class labels flipped. Rotated XOR (R-XOR) rotates XOR by $\theta^{\circ}$ degrees.
This simulation setup facilitates the manipulation of task overlap, allowing for an examination of the transfer properties of our proposed approach under different levels of task similarity (see main text).
A-E Real Data Extended Experiments and Details
This section contains extended results on algorithms not shown in the main text (see Appendix Figure 5). FOOD1k and Mini-Imagenet datasets were obtained from https://www.kaggle.com/datasets/whitemoon/miniimagenet and https://github.com/pranshu28/TAG, respectively.
TABLE III: Benchmark dataset details.
Experiment Dataset Training samples Testing samples Dimension
CIFAR 10X10 CIFAR 100 500050005000 100001000010000 3×32×323$\times$ 32$\times$ 323 × 32 × 32
5-dataset CIFAR-10 500005000050000 100001000010000 3×32×323$\times$ 32$\times$ 323 × 32 × 32 (resized)
MNIST 600006000060000 100001000010000
SVHN 732577325773257 260322603226032
notMNSIT 168531685316853 187318731873
Fashion-MNIST 600006000060000 100001000010000
Split Mini-Imagenet Mini-Imagenet 480004800048000 120001200012000 3×84×843$\times$ 84$\times$ 843 × 84 × 84
FOOD1k 50X20 Food1k 600006000060000 996829968299682 3×50×503$\times$ 50$\times$ 503 × 50 × 50 (resized)
Spoken Digit Spoken Digit 165016501650 135013501350 $\times 28$ 28 (processed and resized)
Split Mini-Imagenet
In this experiment, we have used the Mini-Imagenet dataset [67]. The dataset was split into 202020 tasks with 555 classes each. Each task has 240024002400 training samples and 600600600 testing samples. As shown in Figure 5 right column, we get positive forward and backward transfer for SiLLy-N. However, although samples per task is lower compared to that of 5-dataset, it is still quite high. Hence, Model Zoo outperforms all the algorithms in this experiment.
5-dataset
In this experiment, we have used 5-dataset [67]. It consists of 555 tasks from five different datasets: CIFAR-10 [69], MNIST, SVHN [78], notMNIST [79], Fashion-MNIST [80]. All the monochromatic images were converted to RGB format, and then resized to 3×32×323$\times$ 32$\times$ 323 × 32 × 32. As shown in Appendix Table III, training samples per task in 555-dataset is relatively higher than that of low data regime typically considered in lifelong learning setting. However, as shown in Figure 5 left column, SiLLy-N show less forgetting than most of the reference algorithms. On the other hand, Model Zoo shows comparatively better performance in relatively high task data size setup. Recall that SiLLy-N is based on bagging, and Model Zoo is based on boosting. It is well known that boosting often outperforms bagging when sample sizes are large  [^2] Authors in [81] shows that both bagging and boosting asymptotically converge to the Bayes optimal solution. However, for finite sample size and similar model complexity, we empirically find bagging approach to lifelong learning performs better than that of boosting when the training sample size is low whereas boosting performs better on large training sample size (See Figure 1). This is consistent with similar results in single task learning [82, 83, 43].
Overlapping Task Experiment
We considered the setting where each task is defined by a random sampling of 10 out of 100 classes with replacement in CIFAR 10x10. This environment is designed to demonstrate the effect of having overlapping tasks, which is a common property of real world lifelong learning tasks. Appendix Figure 4 shows positive transfer from other tasks to Task 1 for SiLLy-F and SiLLy-N.
Figure 4: SiLLy-N and SiLLy-F transfer knowledge effectively when tasks share common classes. Each task is a random selection of 10 out of the 100 CIFAR-100 classes. Both SiLLy-F and SiLLy-N demonstrate monotonically increasing transfer efficiency for up to 20 tasks.
Figure 5: Performance of different lifelong learners on two vision datasets..
Spoken Digit Experiment Details
[…]
Figure 6: Spectrogram extracted from eight different recordings of six speakers uttering the digit ‘five’.
Figure 7: Extended results on the Spoken Digit experiments. This plot contains algorithms not shown in main text Figure 1.